A lot of 99 semiconductor chips contains 19 that are defective. (a) Two are selected, one at a time and without replacement from the lot. Determine the probability that the second one is defective. (b) Three are selected, one at a time and without replacement. Find the probability that the first one is defective and the third one is not defective.
Answer #1

(1)

When 1st is defective, 2nd is defective, then
P1 = (19/99)*(18/98)
if the 1st is non-defective, 2nd is defective, then
P2 = (80/99)*(19/98)
so the probability of 2nd being defective
P = P1+P2 = (19/99)*(18/98) + (80/99)*(19/98)

= .1919

(2)de Fective when Si Cond 19 0.029073 97 -0.156 6 6 9 10-1567